# Design of AI-Powered Tool for Self-Regulation Support in Programming Education

### v1

@@ Abstract.
Large Language Model (LLM) tools have demonstrated their potential to deliver high-quality assistance by providing instant, personalized feedback that is crucial for effective programming education. However, many of these tools operate independently from institutional Learning Management Systems, which creates a significant disconnect. This isolation limits the ability to leverage learning materials and exercise context for generating tailored, context-aware feedback. Furthermore, previous research on self-regulated learning and LLM support mainly focused on knowledge acquisition, not the development of important self-regulation skills. To address these challenges, we developed CodeRunner Agent, an LLM-based programming assistant that integrates the CodeRunner, a student-submitted code executing and automated grading plugin in Moodle. CodeRunner Agent empowers educators to customize AI-generated feedback by incorporating detailed context from lecture materials, programming questions, student answers, and execution results. Additionally, it enhances students’ self-regulated learning by providing strategy-based AI responses. This integrated, context-aware, and skill-focused approach offers promising avenues for data-driven improvements in programming education.
LLM-powered tools, Programming education, Self-regulated learning, Personalized feedback, CodeRunner
@@ 1. Introduction
As programming education attracts an ever-growing number of students, providing immediate, individualized assistance during challenging moments has become increasingly difficult (Chiu, 2024; Yilmaz and Yilmaz, 2023b; Ma et al., 2024a). However, the reality in many educational institutions is far from ideal. Traditional support methods, such as scheduled office hours, are often limited, and in-person help can be both time-consuming and labor-intensive. This situation frequently impedes the delivery of the tailored feedback students need to fully comprehend complex programming concepts. Consequently, there is an urgent need for scalable, equitable, and user-friendly solutions that can offer instant support in programming education (Tian et al., [n. d.]; Yilmaz and Yilmaz, 2023a; Pankiewicz and Baker, 2023). In this context, the advent of LLMs presents a promising alternative. A variety of LLM-powered programming assistants have been developed to provide timely coding guidance and suggestions, potentially transforming the landscape of programming education.
@@
While these advancements present exciting opportunities for personalized learning and efficient problem-solving, they also raise critical questions about the role and effectiveness of LLM-powered assistants in truly enhancing student learning. One major concern is that students may become overly reliant on these tools, potentially hindering the development of self-regulated learning (SRL) skills and problem-solving skills (Anagnostopoulos, 2023; Rajala et al., 2023; Humble et al., 2023). Research indicates that the ease of obtaining direct answers can foster a superficial understanding, as students may bypass the rigorous process of working through challenges independently (Skjuve et al., 2023; Tlili et al., 2023). Unfortunately, most existing LLM-powered programming tools fail to address this issue and ignore the implementation of pedagogically sound guardrails to restrict the LLM’s capacity to guide students to engage in self-regulated learning and use AI as a supportive aid (Sun et al., 2024).
@@
Another significant issue is that many LLM-based tools operate independently of institutional Learning Management Systems (LMS) like Moodle. This separation creates a disconnect between the tool and the broader educational context, limiting the opportunity to leverage rich assignment details and course materials for more tailored, context-aware feedback. For example, Ma et al. (Ma et al., 2024a) found that when LLMs are not aligned with a student’s specific curriculum, they can generate advanced or off-topic responses that stray from the intended teaching scope, ultimately detracting from the learning process. Without seamless integration, educators struggle to track how students interact with LLM-generated feedback over time, making it difficult to assess the true impact on learning outcomes (Ma et al., 2024b). Therefore, incorporating LLM-based tools within the LMS environment is essential to ensure that the feedback provided is both relevant and aligned with course objectives. Additionally, the effectiveness of AI-generated feedback is highly sensitive to the context of the learning tasks. The lecture materials and exercises in programming education are highly important contextual information; however, they are not well connected in current LLM-based scaffolding research (Kazemitabaar et al., 2024). The Learning Analytics (LA) technique has to enhance the quality of LLM-based feedback by utilizing the behavioral data in context between learners, lecture materials, and exercise solving (Kaur and Chahal, 2024).
@@
To address these limitations, we developed CodeRunner Agent, a LLM-based programming assistant that seamlessly extends CodeRunner, a free and open-source plug-in for Moodle designed to execute and assess student-submitted code. CodeRunner Agent is designed to meet the needs of both students and educators. It allowed to leverage the comprehensive context available within an LMS environment from learning logs and enhance students’ self-regulated learning in introductory programming education. This enables the delivery of customized feedback that not just give the correct answer but enhances students’ self-regulation and problem solving skills. Beyond timely and strategy-based feedback, the CodeRunner Agent continuously tracks AI responses and subsequent modifications in students’ code. This capability provides educators and researchers with valuable insights into learning trajectories and the overall effectiveness of AI-assisted instruction. By collecting and analyzing data on how students interact with and adapt to AI-generated feedback, our approach paves the way for data-driven improvements in programming education. In summary, CodeRunner Agent represents a significant step forward in integrating AI tools within institutional LMS environments to enhance programming education. By addressing the challenges of scalability, contextual feedback, and skill enhancement, our work offers promising avenues for improving student achievements and deepening our understanding of AI’s role in modern education.
@@ 2. Related Work
Recent studies have primarily focused on assessing LLMs’ capabilities in programming tasks, ranging from code generation and program repair to code explanation and code summarization (Rajala et al., 2023). For instance, Finnie-Ansley et al. (Finnie-Ansley et al., 2023) demonstrated that OpenAI Codex outperforms most students on code-writing questions in both CS1 and CS2 exams. In a similar vein, Savelka et al. (Savelka et al., 2023) evaluated GPT-3 and GPT-4 on programming exercises across three Python courses, revealing that these models progressed from failing typical assessments to passing courses without human intervention. Furthermore, Sarsa et al. (Sarsa et al., 2022) examined programming exercises generated by OpenAI Codex, assessing their novelty, plausibility, and readiness, and highlighted the potential for these models to create effective coding assignments. Recent work by Phung et al. (Phung et al., 2023) systematically compared GPT models with human tutors, finding that they approach human-level performance in both Python programming tasks and the resolution of real-world buggy programs. Additionally, ChatGPT has proven effective in providing feedback on programming assignments and aiding students in applying theoretical knowledge practically. Prior research has underscored the model’s capacity to generate personalized feedback that students rate positively (Pankiewicz and Baker, 2023).
@@
Despite the impressive achievements of these tools in enhancing programming education, they often operate in isolation, lacking the integration of various functionalities and failing to connect with Integrated Development Environments (IDE) and Learning Management Systems (LMS). This fragmented approach does not adequately meet the needs of educators and students, thereby hindering widespread adoption and scalability. To address these limitations, we developed CodeRunner Agent, a comprehensive solution that seamlessly integrates LLM-powered assistance with both IDEs and LMS platforms. By unifying these functionalities, our system offers a more cohesive and effective learning environment, ensuring that both instructors and learners have access to timely, context-aware support on a large scale.
@@ 3.1. Overview of CodeRunner Agent
The framework of CodeRunner Agent is shown in Figure 1. It is developed and integrated to scaffold programming education with an LLM-based assistant in the Moodle LMS. The framework contains a lecture viewer, a CodeRunner plugin and CodeRunner Agent.
@@
CodeRunner Agent is an AI-powered tool to support learners’ SRL and teachers’ customized designs for LLM-based feedback in programming education. It can be embedded into the Moodle LMS and executed with the CodeRunner plugin. Learners can receive general-purpose and programming-specific regulatory strategy hints from LLMs-based feedback. Instructors can upload learning materials to the context engine of the CodeRunner Agent and customize the contextual parameters of the agent to create tailored instruction. The interactions with the CodeRunner Agent are automatically tracked as xAPI statements and stored in the LRS. For instance, the request type, request time, and exercise ID related to the request will be recorded in LRS if learners send a hint request.
@@ 3.2. SRL Support Model in CodeRunner Agent
The SRL scaffoldings in CodeRunner Agent are implemented using a five-phase cycle model, named PPESS (Planning – Program creation – Error correction - Self-monitoring - Self-reflection). The PPESS model is grounded in the well-known Zimmerman’s SRL theory (Zimmerman, 2008) and adopts a conceptual framework for regulating learning in programming (Silva et al., 2024). The model contains five phases: planning, program creation, error correction, self-monitoring, and self-reflection.
The planning phase constitutes the initial stage in code development, encompassing methodologies employed to develop a comprehensive understanding of exercise requirements and identify essential programming elements. The program creation phase represents an implementation stage, wherein learners execute preliminary plans regarding necessary programming components. This phase involves the combination of declarative and procedural knowledge within the programming task, transforming theoretical constructs into functional code. The error correction phase encompasses multiple strategic approaches for addressing coding errors. These include the externalization of meta-cognitive processes to remedy misunderstanding and reflect on limitations, utilization of cognitive strategies for problem identification and solution formulation, and implementation of behavioral regulation techniques. The progress monitoring phase occurs concurrently with the program creation and error correction phases. During this phase, students continuously monitor their progress in completing the task requirements and assess the adequacy of their code regarding programming syntax and standards. Finally, the self-reflection phase, occurring at the end of code development, enables learners to identify areas requiring improvement, maintain motivation throughout their programming, and engage in evaluation regarding performance and behavior for enhancing learning experiences and skill development.
The phase, target SRL strategy, and LLMs support in the CodeRunner Agent are summarized in Table 1. Both general-purpose and programming-specific regulatory strategies are included for SRL scaffoldings in the five-phase PPESS model. To enhance strategy-based programming learning for novice learners, LLM-powered feedback is integrated into the five-phase PPESS model by leveraging LLMs’ capabilities as personalized, adaptive scaffolds.
Table 1. Phase, Target SRL strategy, LLMs support in the CodeRunner Agent.
@@
Planning Problem understanding, problem definition, program logic planning LLMs offer the basic knowledge of the exercise’s requirements and suggest planning the program step-by-step using diagrams, pseudocode, or notes.
Program creation Review lecture materials, review previous exercise, code dividing, code commenting LLMs provide the location of required knowledge in lecture materials, supplemental resources related to the exercise, and explanations for the key points.
Error correction Review the exercise statement, utilize test cases, analyze the error message, emotional regulation, help-seeking LLMs provide suggestions for effective error correction and generate hints on fixing syntactic and logical errors without showing the solution directly.
Self-monitoring Check the progress of all exercises, test the program regularly LLMs encourage learners to track their own learning progress regularly.
Self-reflection Achievement self-assessment, effort self-assessment, code review, code refactoring LLMs give evaluations on students’ behavioral process and final performance, motivate students by finding their strength points or the effort they put in, and suggest students to identify their improvement areas.
@@
The user interface of the integrated CodeRunner Agent for learners is shown in Figure 2. The dialogue window of the CodeRunner Agent is displayed immediately below the test results of CodeRunner. Learners can select one SRL phase from five phases: planning, program creation, error correction, self-monitoring, and self-reflection. They can choose one request type from two types: general purpose or programming-specific. After sending the hint request, the LLM-based strategy hint will be shown in the response window of the CodeRunner Agent.
@@ 3.3. Context Engine in CodeRunner Agent
The CodeRunner Agent’s core intelligence contains two context engines: the Learning Analytics Context Engine (LACE) and the Knowledge Context Engine (KCE). The LACE calculates and aggregates learners’ engagement metrics (i.e., time spent, attempt frequency) and performance metrics (i.e., success rates, error patterns) from LRS data after filtering personal data (i.e., name, gender, and email). The KCE handles both lecture and exercise knowledge bases. The lecture knowledge base manages the key programming concepts and the dependencies between the concepts from lecture materials. On the other hand, the exercise knowledge base categorizes exercises by concept, difficulty, solution, and typical mistakes.
Instructors can upload lecture materials (i.e., concept definition, concept illustration, and annotated examples) and exercises (i.e., problem statement, solution, and test cases) to the context engine. The lecture materials and exercises will be converted to textual knowledge and stored in the knowledge database for future retrieval. More importantly, instructors can update the knowledge bases and customize the parameters of the context engine by using the configuration tools.
@@ 4. Discussion and Future work
This study aims to develop an LLM-based programming assistant, CodeRunner Agent, that seamlessly integrates with a lecture viewer and CodeRunner plugin in the Moodle LMS. The design is grounded within Zimmerman’s SRL theory and targets the freshmen students’ programming-specific knowledge acquisition as well as general self-regulation skill development. This study can fill the gap between AI, SRL, and learning analytics (LA) by combining LLM-based scaffolding, self-regulation in programming, and learning log-based contextual feedback.

### v2

@@ Abstract.
Large Language Model (LLM) tools have demonstrated their potential to deliver high-quality assistance by providing instant, personalized feedback that is crucial for effective programming education. However, many of these tools operate independently from institutional learning management systems, which creates a significant disconnect. This isolation limits the ability to leverage learning material and exercise contexts for generating tailored, context-aware feedback. Furthermore, previous research on LLM support for programming learning mainly focused on knowledge acquisition, not the development of important self-regulation skills. To address these challenges, we designed CodeRunner Agent, an LLM-based programming tool that integrates the CodeRunner, a student-submitted code executing and automated grading plugin in Moodle. CodeRunner Agent enhances students’ contextual self-regulated learning by providing learning log-based contextual feedback and self-regulation strategy-based AI scaffolding. Additionally, CodeRunner Agent empowers educators to customize AI-generated feedback by incorporating detailed context from lecture materials, programming questions, student answers, and execution results. This integrated, context-aware, and skill-focused approach offers a promising avenue for data-driven programming education.
LLM-powered tool, Self-regulated learning, Programming education, Personalized feedback, Learning analytics
@@ 1. Introduction
Programming has become an increasingly important part of university education; however, it is becoming more challenging for educators to provide timely, personalized support to each student(Chiu, 2024; Yilmaz and Yilmaz, 2023b; Ma et al., 2024a). Traditional support methods, such as scheduled office hours, are often limited, and in-person help can be both time-consuming and labor-intensive. In this context, Large Language Models (LLMs) are emerging as a promising solution to this challenge, offering the potential for on-demand, personalized programming support that can supplement traditional teaching methods(Yilmaz and Yilmaz, 2023a; PANKIEWICZ and BAKER, 2023). A variety of LLM-powered programming assistants have been developed to provide timely coding guidance and suggestions, potentially transforming the landscape of programming education.
@@
While these advancements present exciting opportunities for personalized learning and efficient problem-solving, they also raise critical questions about the role and effectiveness of LLM-powered assistants in truly enhancing student learning. One major concern is that students may become overly reliant on these tools, potentially hindering the development of self-regulated learning (SRL) skills and problem-solving skills (Gong et al., 2024; Rajala et al., 2023; Humble et al., 2023). Research suggests that the convenience of receiving direct answers from LLMs may hinder the development of SRL skills, as students may avoid the deeper cognitive effort required to work through challenges independently(Prasad and Sane, 2024; Skjuve et al., 2023; Tlili et al., 2023). This harmful effect becomes more serious in the introductory programming in university since it’s a challenging process in programming learning for freshman students. Unfortunately, most existing LLM-powered programming tools fail to address this issue and ignore the implementation of pedagogically sound scaffolding to enhance students’ self-regulated learning while avoiding LLMs’ harmful effects(Becker et al., 2023; Sun et al., 2024).
@@
Another significant issue is that many LLM-based tools operate independently of institutional Learning Management Systems (LMS) like Moodle. This separation creates a disconnect between the tool and the broader educational context such as course materials and rich assignment details. For example, Ma et al. (Ma et al., 2024a) found that when LLMs are not aligned with a student’s specific curriculum, they generate advanced or off-topic responses that stray from the intended teaching scope, ultimately detracting from the learning process. Without seamless integration, educators struggle to track how students interact with LLM-generated feedback over time, making it difficult to assess the true impact on learning outcomes (Ma et al., 2024b). Therefore, it’s essential to ensure that the AI-generated feedback is both relevant and aligned with course objectives by incorporating LLM-based tools within the LMS environment. The lecture materials and exercises in programming education are highly important contextual information; however, they are not well connected in current LLM-based scaffolding research (Kazemitabaar et al., 2024). The Learning Analytics (LA) technique can enhance the quality of LLM-based feedback by utilizing the behavioral data in context between learners, lecture materials, and exercise solving (Kaur and Chahal, 2024).
@@
To address these limitations, we designed CodeRunner Agent, a LLM-based programming tool that seamlessly extends CodeRunner, a free and open-source plug-in in Moodle designed to execute and assess student-submitted code. CodeRunner Agent is designed to meet the needs of both learners and educators. It leverages the comprehensive context available within an LMS environment from learning logs and enhances students’ self-regulated learning in introductory programming education. Specifically, it enhances the delivery of individualized feedback by combining students’ knowledge level, self-regulated behaviors, and strategy-focused LLM-based scaffoldings. Beyond contextual and strategy-based feedback, CodeRunner Agent tracks students’ requests and AI responses. This capability provides educators and researchers with valuable insights into AI-powered learning process and the overall effectiveness of AI-assisted instruction. By collecting and analyzing data on how students interact with AI-generated feedback, our approach paves the way for data-driven improvements in programming education. In summary, our proposal represents a significant step forward in integrating AI tools within institutional LMS environments to enhance programming education. By addressing the challenges of AI integration, contextual feedback, and self-regulation scaffolding, our work offers promising avenues for enhancing student skill development and deepening our understanding of AI’s role in modern education.
@@ 2. Related Work
Recent studies have primarily focused on assessing LLMs’ capabilities in programming tasks, ranging from code generation and program repair to code explanation and code summarization (Rajala et al., 2023; Chen et al., 2023). For instance, Finnie-Ansley et al. (Finnie-Ansley et al., 2023) demonstrated that OpenAI Codex outperforms most students on code-writing questions in both CS1 and CS2 exams. In a similar vein, Savelka et al. (Savelka et al., 2023) evaluated GPT-3 and GPT-4 on programming exercises across three Python courses, revealing that these models progressed from failing typical assessments to passing courses without human intervention. Furthermore, Sarsa et al. (Sarsa et al., 2022) examined programming exercises generated by OpenAI Codex, assessing their novelty, plausibility, and readiness, and highlighted the potential for these models to create effective coding assignments. Recent work by Phung et al. (Phung et al., 2023) systematically compared GPT models with human tutors, finding that they approach human-level performance in both Python programming tasks and the resolution of real-world buggy programs. Additionally, ChatGPT has proven effective in providing feedback on programming assignments and aiding students in applying theoretical knowledge practically. Prior research has underscored the model’s capacity to generate personalized feedback that students rate positively (PANKIEWICZ and BAKER, 2023).
@@
Despite the impressive achievements of these tools in enhancing programming education, they often operate in isolation, lacking the integration of various functionalities and failing to connect with existed Learning Management Systems (LMS). This fragmented approach does not adequately meet the needs of educators and students, thereby hindering widespread adoption and scalability. To address these limitations, we developed CodeRunner Agent, a comprehensive solution that seamlessly integrates LLM-powered assistance with LMS platforms. By unifying these functionalities, our system offers a more cohesive and effective learning environment, ensuring that both instructors and learners have access to timely, context-aware support on a large scale.
@@ 3.1. Overview of CodeRunner Agent
The framework of CodeRunner Agent is shown in Figure 1. It is designed and developed to scaffold programming education with an LLM-based tool in the Moodle LMS. The framework contains a lecture viewer, a CodeRunner plugin and CodeRunner Agent.
@@
CodeRunner Agent is an AI-powered tool to support learners’ SRL and teachers’ customized designs for LLM-based feedback in programming education. It can be embedded into the Moodle LMS and executed with the CodeRunner plugin. Learners can receive general-purpose and programming-specific regulatory strategy hints from LLMs-based feedback. Instructors can upload learning materials to the context engine of the CodeRunner Agent and customize the contextual parameters of the agent to create tailored instruction. The interactions with the CodeRunner Agent are automatically tracked as xAPI statements and stored in the LRS. For instance, the request type, request time, and exercise ID related to the request will be recorded in LRS if learners send a request to the agent.
@@ 3.2. SRL Support Model in CodeRunner Agent
The SRL scaffoldings in CodeRunner Agent are implemented using a five-phase cycle model named PPESS: Planning, Program creation, Error correction, Self-monitoring, Self-reflection. The PPESS model is grounded in the well-known Zimmerman’s SRL theory (Zimmerman, 2008) and specified for self-regulation in programming learning (Silva et al., 2024).
The phase, target SRL strategy, and LLMs support in the CodeRunner Agent are summarized in Table 1. Both general-purpose and programming-specific regulatory strategies are included for SRL scaffoldings in the five-phase PPESS model.
The Planning phase serves as the foundational stage in programming, including cognitive and metacognitive strategies for problem understanding, problem decomposition, solution architecture and essential programming component identification. The Program creation phase represents an implementation stage, wherein learners execute preliminary plans by implementing programming constructs that address identified requirements. This phase involves the combination of declarative and procedural knowledge within the programming task, facilitating the transfer of coding knowledge. The Error correction phase encompasses sophisticated learning strategies for addressing coding errors. These include the externalization of metacognitive processes to remediate conceptual misunderstandings, employment of cognitive diagnosis strategies for systematic problem identification, and implementation of self-regulation to optimize the debugging process. The Self-monitoring phase occurs concurrently with the program creation and error correction phases, representing an ongoing metacognitive process. Learners engage in continuously assessment and adjustment toward task completion and code quality. Finally, the Self-reflection phase, occurring at the end of programming, enables learners to comprehensively evaluate their programming experience by identifying areas requiring improvement, maintaining motivation, and improving current learning strategies for future tasks.
Table 1. Phase, Target SRL strategy, LLM-powered feedback in the CodeRunner Agent.
@@
Planning Problem understanding, problem definition, program logic planning Offer the basic knowledge of the exercise’s requirements and suggest planning the program step-by-step using diagrams, pseudocode, or notes.
Program creation Review lecture materials, review previous exercises, code dividing, code commenting Provide the location of required knowledge in lecture materials, supplemental resources related to the exercise, and explanations for the key points.
Error correction Review the exercise statement, utilize test cases, analyze the error message, help-seeking Give suggestions for effective error correction and generate hints on fixing syntactic and logical errors without showing the solution directly.
Self-monitoring Check exercises progress, test the program regularly Encourage learners to track their own learning progress regularly.
Self-reflection Achievement self-assessment, effort self-assessment, code review, code optimization Provide evaluations on learners’ behavioral process and final performance, motivate learners by finding their strength points or the effort they put in, and suggest learners to identify their improvement areas.
The user interface of the integrated CodeRunner Agent for learners is shown in Figure 2. It contains three main components: (a) Question & Answer, (b) Check with Test Cases, and (c) LLM-based Support. Firstly, learners check the question statement for programming exercise and input their code solutions for the question in the top Question & Answer component. Secondly, learners submit their code solutions for checking pre-defined test cases by comparing the execution results with the expected outputs in the central Check with Test Cases component. Finally, learners have the option to select one of five SRL phases like error correction, choose one request type from general purpose or programming-specific, and send the request to the agent in the bottom LLM-based Support. Upon submitting a request, the LLM-based strategy-focused feedback is displayed within the response box.
@@ 3.3. Context Engine in CodeRunner Agent
The CodeRunner Agent’s core intelligence contains two context engines: Learning Analytics Context Engine (LACE) and Knowledge Context Engine (KCE). LACE calculates and aggregates learners’ engagement metrics (i.e., time spent, attempt frequency) and performance metrics (i.e., success rates, error patterns) from LRS data after filtering personal data (i.e., name, gender, and email). LACE is integrated as contextual embedding of self-regulation behavioral strategies with the feedback of the agent. KCE handles both lecture and exercise knowledge bases. The lecture knowledge base manages the key programming concepts and the dependencies between the concepts from lecture materials. The exercise knowledge base categorizes exercises by concept, difficulty, solution, and typical mistakes. KCE is used as learner knowledge embedding in the agent.
Instructors can upload lecture materials (i.e., concept definition, concept illustration, and annotated examples) and exercises (i.e., problem statement, solution, and test cases) to the context engine. They will be converted to textual knowledge and stored in the knowledge database for future retrieval. More importantly, instructors can update the knowledge bases and customize the parameters of the context engine by using the configuration tools.
@@ 4. Discussion and Future work
This study aims to design an LLM-based programming tool, CodeRunner Agent, that seamlessly integrates with a lecture viewer and CodeRunner plugin in the Moodle LMS. The design is grounded within Zimmerman’s SRL theory and targets the freshmen students’ programming-specific knowledge acquisition as well as general self-regulation skill development. The agent provides learning log-based contextual feedback and self-regulation strategy-based AI scaffolding to enhance contextual SRL. Furthermore, the agent empowers educators to customize AI-generated feedback by incorporating detailed context from lecture materials, programming questions, student answers, and execution results. This study can fill the gap between AI, SRL, and LA by combining learning log-based contextual feedback, LLM-based self-regulation scaffolding, and seamlessly integrating in LMS.